OpenStudy (anonymous):

Evaluate the expression 5(7 − 4)2 ÷ 3 + 11.

OpenStudy (anonymous):

that 2 is an exponent by the way

OpenStudy (anonymous):

Use PEMDAS.

pooja195 (pooja195):

Use PEMDAS Parentheses Exponents Multiplication Division Addition Subtraction

pooja195 (pooja195):

Start with 7-4 = ?

OpenStudy (anonymous):

3

pooja195 (pooja195):

good so now we have \[\huge~\rm~\frac{ 5(3^2) }{ 3 }+11\] Whats 3^2=?

pooja195 (pooja195):

the fraction just means divided by 3

OpenStudy (anonymous):

21

OpenStudy (anonymous):

That is with using distributive property

OpenStudy (anonymous):

5(7 − 4)2 ÷ 3 + 11. (5)(3)(2) /3 +11 (15) (2) / 3 + 11 30/3 + 11 10 + 11 = 21